# Algorithmic Behaviors Across Regions: A Geolocation Audit of YouTube Search for COVID-19 Misinformation between the United States and South Africa

### v1

@@ Appendix A Model Fine-tuning
We selected the Sentence Transformer model all-mpnet-base-v2 due to their best performance generating sentence embeddings across various tasks (Song et al. 2020). However, all-mpnet-base-v2 was not trained on COVID-19-related texts. Scialom et al. (2022) demonstrated that language models can be “continual learners” of new domains via self-supervised training. Thus, we continued training the model using a self-supervised contrastive learning framework (i.e. multiple-negative-ranking-loss) on COVID-19 question-answering (QA) datasets (Möller et al. 2020; Lu 2022; Tang et al. 2020), totaling 3,296 QA pairs, to generate more meaningful sentence embeddings on the COVID-19-related texts. We randomly partitioned the dataset into 80% training, 10% validation, and 10% testing sets. To avoid overfitting, we trained the model until the training loss decreased further relative to the validation loss. Our best model used an Adam optimizer with a learning rate of 2e-5, 3 epochs, 32 batch size, and a warmup step for 10% of total steps. We found that our fine-tuning worked. The finetuned model had a test loss of 0.7466 and the original pre-trained model had a test loss of 1.7722, indicating that the finetuned model generated more meaningful sentence embeddings on COVID-19 pandemic-related texts.
@@ Appendix B Measuring Geolocation-based Personalization
To measure geolocation-based personalization, we used the Jaccard index to calculate the similarity of two SERPs. Previous audits used the Jaccard index to measure personalization in web searches (Kliman-Silver et al. 2015; Juneja et al. 2021). The Jaccard index determines the similarity between two lists – 1 indicates that both lists have identical elements, while 0 indicates that both have completely different elements.
@@ Appendix E Validating Proxies for Accurate IP Location
After obtaining unsuccessful validation results from the geospoofing method, we conducted a third experiment to validate that IPRoyal provides the bots with accurate IP geolocations. We created 2 sets of twin bots in Cape Town, South Africa on AWS, giving us 4 bots total. A set of twin bots would search regularly, while the other set of twin bots would search using the IPRoyal proxies set to Cape Town, South Africa. Since we are comparing bots in the same geolocation, we should ideally see no geolocation-based personalization. Additionally, we used IP2Location to keep track of IP addresses (IP2Location 2023), enabling us to validate whether the IPRoyal proxies are associated with our desired IP geolocation. Since the 2 sets of twin bots were at the same IP geolocation, our GBP metric value should be close to 0. From December 22nd to December 24th, 2022, we ran the validation experiment at two evenly distributed times of 00:00 UTC and 12:00 UTC. The average geolocation-based personalization value was -0.02171717Note that, our formalized metric from Appendix B may result in a slight negative value since we take the more restrictive noise to establish our baseline noise. Regardless, any slightly negative value indicates that the differences between the treatments did not exceed the baseline noise., validating that the bots had no geolocation-based personalization. Additionally, our IP2Location data also validates the consistency and accuracy of IPRoyal in providing proxies with valid IP geolocations in Cape Town.
Name of Proxy Provider Price (\$) Proxy Geolocations in South Africa (City)
@@
IPRoyal \$2-4 / GB of traffic Cape Town, Johannesburg, Durban, Pretoria
@@
Table 6: Proxy providers with services available in South Africa. We list each proxy provider’s price (per gigabyte of traffic data or monthly subscription) and the proxy geolocations available in South Africa. Some proxy providers do not have a dedicated pool of proxies in specific geolocations/cities in South Africa, which we denoted as “N/A” in the column “Proxy Geolocations in South Africa (City).” Except for IPRoyal, all proxy providers were expensive (e.g. $>$\$10 / GB of traffic) or did not have proxies available in our selected geolocations in SA.
Scale Annotation Label Descriptions Example Video (Video Title and ID)
-1 Opposing COVID-19 Misinformation (Opposing) The video opposes, debunks, satirizes OR provides countervailing contexts to COVID-19 misinformation OR disseminates/endorses public health authorities’ policies OR explains the scientific contexts surrounding the public health policy and the COVID-19 pandemic OR provides countervailing contexts to the usage of misleading terms, such as “Kungflu” and “China Virus”. 1. Fox Allows Lunatic to Imply Kim Jong-Un Created Cronavirus (Video ID: 8EGVzWRic1I)
2. The Monday Times: Coronavirus Car, Ghosts, Tik Tok, 5G Conspiracy — The Daily Show With Trevor Noah (Video ID: 2totKKH4770)
0 Neutral COVID-19 Information (Neutral) Videos that broadly cover anything related to the COVID-19 pandemic, but do not support nor oppose COVID-19 misinformation. For example, videos that report on the number of COVID-19 cases, discuss remote work hours during a pandemic, etc. 1. Coronavirus outbreak: Hospitals in China swamped with patients (Video ID: 9qAFanPeKtE)
2. Africa Unite Against COVID 19 (Video ID: 6EMEEn65JPg)
1 Supporting COVID-19 Misinformation (Supporting) The video supports OR provides evidence for COVID-19 misinformation OR suggests alternative treatment, prevention, and cure without scientific evidence OR exaggerates a claim to suggest a misinformative narrative OR promotes the usage of misleading terms without countervailing contexts. 1. KT The Arch Degree Break’s Down The Coronavirus! (MUST SEE) (Video ID: ulncPbTgPlo)
2. Origin of the CCP Virus: Fauci Emails Revealed; Ivermectin: A Vaccine Alternative?(Video ID: edT1HYMjnzA)
2 On the COVID-19 origins in Wuhan, China (COVID-19 origins) Any videos that hypothesize OR cover the origins of COVID-19 in Wuhan, China as the main premise of the video (e.g. natural origins theory, lab leak theory) OR provide evidence for the lab leak theory OR trying to debunk/satirize the lab leak theory. Videos that speculate beyond the origin theories of the virus (e.g., COVID-19 was intentionally engineered as a bioweapon in the Wuhan Lab) are marked as supporting COVID-19 misinformation. 1. Progressive Lab Leak Theory (Video ID: 0iNZE6BEoXs)
2. Why the US is taking a second look at the ’lab leak’ theory about COVID-19 (Video ID: ZRV_uYS2Buc)
3 Irrelevant Any video whose content is not related to the COVID-19 pandemic and vaccines. 1. Dengue: Laboratory-acquired case reported in North Carolina, according to study (Video ID: PxFJ3YSgG0Y)
2. The Nuremberg Code (Video ID: Eu_3Cbf7f8Y)
4 YouTube video in a language other than English (Non-English) The video’s title, description, and/or contents in a language other than English. 1. Can Coronavirus be Biological Weapon? Video Analysis by Major General S B Asthana (Retd) (Video ID: WiHmN_DuQhQ)
2. Shiva $\mid$ The Arrogant Kung Fu Fighter $\mid$ Episode 54 $\mid$ Download Voot Kids App (Video ID: ISfm5pAWmxs)
@@ Appendix F Amazon Mechanical Turk Job
We took two steps to ensure the instructions, test questions, and overall annotation task were clear and comprehensive. First, we posted the Qualification Test and the annotation task on r/mturk [^18] https://www.reddit.com/r/mturk/, a subreddit community of AMT workers, and Turker Nation, an unofficial Slack channel of AMT workers. Second, we conducted a pilot run by posting 13 annotation tasks and the aforementioned screening requirements. Our task received positive feedback from the AMT community, which indicates acknowledgment of our task design. One respondent reported “Seems interesting. Hope to get to work on it when it’s ready,” while another stated “The test was fairly easy and constructed well.” After obtaining positive feedback from the AMT communities and a successful pilot run, we released our AMT annotation task titled “YouTube Video Labeling (COVID-19).” We paid the AMT workers the United States federal minimum wage (\$7.25 per hour). When refining our annotation heuristics in §4, we kept track of the time it took the external researchers to annotate the 13 videos, which ranged from 26 seconds to 42 minutes in video duration. We found that it took, on average, 1.5 minutes to annotate a video. However, we erred on the side of overestimation and used 2 minutes as the baseline to annotate a video. We prorated the federal minimum wage using the 2-minute annotation time, thus paying each annotator \$0.24 per annotation. Our compensation is much higher than prior works (Papadamou et al. 2022; Albadi et al. 2022; Liaw et al. 2023); for example, Papadamou et al. (2022) paid \$0.03 per annotation for labeling YouTube videos for pseudoscientific content.
@@
In our annotation task, we required each AMT worker to assign an annotation label to the video and provide a short rationale behind selecting the label. We required a minimum of 10 characters for our rationale. We released 1,988 videos on AMT in batches of 15 at a time. Three different AMT workers annotated each video. Throughout the annotation process, we sampled videos in each batch to ensure quality work and reasonable rationales from the AMT workers [^19] We found that two AMT workers submitted identical annotations and rationale across the tasks; we rejected their work.. Using the collected annotations, we calculated the Fleiss’ Kappa Score ($\kappa$) to assess the annotators’ agreement. We obtained $\kappa=0.62$, which is considered “substantial agreement.” Our $\kappa$ value of 0.62 is a much higher agreement score compared to a close prior work; Papadamou et al. (2022) asked three AMT workers to annotate YouTube videos for pseudoscientific content, including COVID-19 conspiracy theories, using 3 labels, resulting in a Fleiss’ $\kappa$ score of 0.14.
@@ Appendix G Classifier Training Details
Video Comments: Serrano et al. (2020) found that YouTube comments was a highly predictive feature for detecting COVID-19 misinformation in YouTube videos. Thus, we selected the top 100 comments associated with each video.
Model Hyperparameter Search Space
SVM
C: [0.1, 1, 10, 100, 1000]
gamma: [1, 0.1, 0.01, 0.001, 0.0001]
Kernel: [rbf]
XGBoost
max_depth: [2, 3, 4, 5, 6, 7, 8, 9, 10]
n_estimators: [60, 100, 140, 180, 220]
learning_rate: [0.5, 0.1, 0.01, 0.05, 0.001]
Table 8: The table displaying the hyperparameter search space employed in each traditional machine learning model. To find the most optimal set of hyperparameters for each model, we employed the grid search strategy, where we exhaustively trained and evaluated the model over all possible hyperparameter combinations in the search space. The grid search strategy was conducted using five-fold cross-validation, searching for a set of hyperparameters with the highest average cross-validation accuracy.
@@
Support Vector Machines and XGBoost
@@
DeBerta-v3-base and DeBerta-v3-large
@@
GPT-4 Turbo
We describe the prompt design considerations and tuning process to evaluate GPT-4 Turbo’s performance in detecting COVID-19 misinformation based on YouTube video metadata. Since we employed an OpenAI’s GPT-4 model, we designed our prompts based on OpenAI’s recommendations on prompt-engineering (OpenAI 2024) and prior work (Mishra and Chatterjee 2023). First, prior research showed that models demonstrate improved performance when compelled to reason and justify their decisions (OpenAI 2024; Dammu et al. 2024). Therefore, we mandated GPT-4 to provide direct excerpts from the YouTube video metadata and concise justifications regarding their selected label. Second, OpenAI (2024) recommended that asking the model to adopt a “persona” in their system can lead to better results from LLMs. Thus, for our task, we prompted GPT-4 to adopt a persona as a “public health expert” with comprehensive knowledge of the COVID-19 pandemic and its misinformation. Third, we experimented with both zero-shot prompting and few-shot prompting. Zero-shot prompts involve presenting the task to the LLM, only including the label descriptions without accompanying examples or training. Meanwhile, few-shot prompts involve including task-specific examples within the prompt, enabling the pre-trained LLM to condition on the illustrative examples rather than updating its weights. We provided five few-shot examples, each containing video metadata, an assigned label, and reasoning for the label. These few-shot examples were the same as the ones provided to AMT workers on their annotation tasks. Fourth, we experimented with varying temperature levels to find our task’s most ideal parameter setting. Ranging from 0 to 2, the temperature parameter influences how models generate text. When utilizing lower values for temperature, such as 0, the generated text becomes deterministic, selecting more consistent and coherent outputs. However, higher values for temperature introduce randomness, generating more diverse and creative results. Previous works (Mishra and Chatterjee 2023) also experimented with various temperature values, such as 0.2 and 0.7, for their text annotation tasks and found that lower temperature values (e.g. 0.2) resulted in the best performance. Likewise, we selected relatively lower temperature values: 0, 0.2, and 0.7 for our experiments. We used default settings for other parameters. See Table 11 for our zero-shot prompt and Table 12 for our few-shot prompt.
@@ Appendix H Classifier Performance Results
Performance results for each individual feature and combination of features are presented in Index 0-13 of Table 10. Among the individual features, we observed that video titles yielded the highest performance (0.64 accuracy), suggesting that they are an informative input feature. Similarly, descriptions (0.6 accuracy) and transcripts (0.64 accuracy) demonstrated comparable performance. In contrast to the findings of Serrano et al. (2020), where they identified video comments as highly predictive in detecting COVID-19 misinformation in YouTube videos, our results revealed video comment as the least effective input feature, achieving an accuracy of only 0.49. Combining title, description, and transcript (Index 9) or title, description, transcript, and tags (Index 12) yielded the best performance of 0.69 accuracy. This further indicates that video comments are not informative for our classification task.

### v2

@@ 10 Conclusion
Acknowledgements
Mitra and Jung were partially supported by National Institute of Health’s NIDA grant DA056725-01A1. Jung was also supported by the Mary Gates Endowment and the National Science Foundation grant #2128642. We thank the Poynter Fact-Checking Innovation Initiative award for supporting this work, PesaCheck for the fact-checked video dataset, and AMT workers for their annotations.
@@ Appendix A Model Fine-tuning
We selected the Sentence Transformer model all-mpnet-base-v2 due to their best performance generating sentence embeddings across various tasks (Song et al. 2020). However, all-mpnet-base-v2 was not trained on COVID-19-related texts. Scialom, Chakrabarty, and Muresan (2022) demonstrated that language models can be “continual learners” of new domains via self-supervised training. Thus, we continued training the model using a self-supervised contrastive learning framework (i.e. multiple-negative-ranking-loss) on COVID-19 question-answering (QA) datasets (Möller et al. 2020; Lu 2022; Tang et al. 2020), totaling 3,296 QA pairs, to generate more meaningful sentence embeddings on the COVID-19-related texts. We randomly partitioned the dataset into 80% training, 10% validation, and 10% testing sets. To avoid overfitting, we trained the model until the training loss decreased further relative to the validation loss. Our best model used an Adam optimizer with a learning rate of 2e-5, 3 epochs, 32 batch size, and a warmup step for 10% of total steps. We found that our fine-tuning worked. The finetuned model had a test loss of 0.7466 and the original pre-trained model had a test loss of 1.7722, indicating that the finetuned model generated more meaningful sentence embeddings on COVID-19 pandemic-related texts.
@@ Appendix B Measuring Geolocation-based Personalization
To measure geolocation-based personalization, we used the Jaccard index to calculate the similarity of two SERPs. Previous audits used the Jaccard index to measure personalization in web searches (Kliman-Silver et al. 2015; Juneja and Mitra 2021). The Jaccard index determines the similarity between two lists – 1 indicates that both lists have identical elements, while 0 indicates that both have completely different elements.
@@ Appendix E Validating Proxies for Accurate IP Location
After obtaining unsuccessful validation results from the geospoofing method, we conducted a third experiment to validate that IPRoyal provides the bots with accurate IP geolocations. We created 2 sets of twin bots in Cape Town, South Africa on AWS, giving us 4 bots total. A set of twin bots would search regularly, while the other set of twin bots would search using the IPRoyal proxies set to Cape Town, South Africa. Since we are comparing bots in the same geolocation, we should ideally see no geolocation-based personalization. Additionally, we used IP2Location to keep track of IP addresses (IP2Location 2023), enabling us to validate whether the IPRoyal proxies are associated with our desired IP geolocation. Since the 2 sets of twin bots were at the same IP geolocation, our GBP metric value should be close to 0. From December 22nd to December 24th, 2022, we ran the validation experiment at two evenly distributed times of 00:00 UTC and 12:00 UTC. The average geolocation-based personalization value was -0.02, [^18] Note that, our formalized metric from Appendix B may result in a slight negative value since we take the more restrictive noise to establish our baseline noise. Regardless, any slightly negative value indicates that the differences between the treatments did not exceed the baseline noise. validating that the bots had no geolocation-based personalization. Additionally, our IP2Location data also validates the consistency and accuracy of IPRoyal in providing proxies with valid IP geolocations in Cape Town.
Proxy Provider Price (\$) Proxy Geolocations in South Africa (City)
@@
IPRoyal \$2–4 / GB of traffic Cape Town, Johannesburg, Durban, Pretoria
@@
Table 6: Proxy providers with services available in South Africa. We list each proxy provider’s price (per gigabyte of traffic data or monthly subscription) and the proxy geolocations available in South Africa. Some proxy providers do not have a dedicated pool of proxies in specific geolocations/cities in South Africa, which we denoted as “N/A” in the column “Proxy Geolocations in South Africa (City).” Except for IPRoyal, all proxy providers were expensive (e.g., $>$\$10 / GB of traffic) or did not have proxies available in our selected geolocations in SA.
@@ Appendix F Amazon Mechanical Turk Job
We took two steps to ensure the instructions, test questions, and overall annotation task were clear and comprehensive. First, we posted the Qualification Test and the annotation task on r/mturk, [^19] https://www.reddit.com/r/mturk/ a subreddit community of AMT workers, and Turker Nation, an unofficial Slack channel of AMT workers. Second, we conducted a pilot run by posting 13 annotation tasks and the aforementioned screening requirements. Our task received positive feedback from the AMT community, which indicates acknowledgment of our task design. One respondent reported “Seems interesting. Hope to get to work on it when it’s ready,” while another stated “The test was fairly easy and constructed well.” After obtaining positive feedback from the AMT communities and a successful pilot run, we released our AMT annotation task titled “YouTube Video Labeling (COVID-19).” We paid the AMT workers the United States federal minimum wage (\$7.25 per hour). When refining our annotation heuristics in §4, we kept track of the time it took the external researchers to annotate the 13 videos, which ranged from 26 seconds to 42 minutes in video duration. We found that it took, on average, 1.5 minutes to annotate a video. However, we erred on the side of overestimation and used 2 minutes as the baseline to annotate a video. We prorated the federal minimum wage using the 2-minute annotation time, thus paying each annotator \$0.24 per annotation. Our compensation is much higher than prior works (Papadamou et al. 2022; Albadi, Kurdi, and Mishra 2022; Yi Liaw et al. 2023); for example, Papadamou et al. (2022) paid \$0.03 per annotation for labeling YouTube videos for pseudoscientific content.
@@
In our annotation task, we required each AMT worker to assign an annotation label to the video and provide a short rationale behind selecting the label. We required a minimum of 10 characters for our rationale. We released 1,988 videos on AMT in batches of 15 at a time. Three different AMT workers annotated each video. Throughout the annotation process, we sampled videos in each batch to ensure quality work and reasonable rationales from the AMT workers. [^20] We found that two AMT workers submitted identical annotations and rationale across the tasks; we rejected their work. Using the collected annotations, we calculated the Fleiss’ Kappa Score ($\kappa$) to assess the annotators’ agreement. We obtained $\kappa=0.62$, which is considered “substantial agreement.” Our $\kappa$ value of 0.62 is a much higher agreement score compared to a close prior work; Papadamou et al. (2022) asked three AMT workers to annotate YouTube videos for pseudoscientific content, including COVID-19 conspiracy theories, using 3 labels, resulting in a Fleiss’ $\kappa$ score of 0.14.
@@ Appendix G Classifier Training Details
Video Comments: Medina Serrano, Papakyriakopoulos, and Hegelich (2020) found that YouTube comments was a highly predictive feature for detecting COVID-19 misinformation in YouTube videos. Thus, we selected the top 100 comments associated with each video.
@@
Model Hyperparameter Search Space
SVM C: [0.1, 1, 10, 100, 1000] $\gamma$: [1, 0.1, 0.01, 0.001, 0.0001] Kernel: [rbf]
XGBoost max_depth: [2, 3, 4, 5, 6, 7, 8, 9, 10] n_estimators: [60, 100, 140, 180, 220] learning_rate: [0.5, 0.1, 0.01, 0.05, 0.001]
Table 7: Hyperparameter search space for each ML model. Grid search with 5-fold cross-validation was used to find optimal hyperparameters with highest average accuracy.
Support Vector Machines and XGBoost.
@@
DeBerta-v3-base and DeBerta-v3-large.
@@
GPT-4 Turbo.
We describe the prompt design considerations and tuning process to evaluate GPT-4 Turbo’s performance in detecting COVID-19 misinformation based on YouTube video metadata. Since we employed an OpenAI’s GPT-4 model, we designed our prompts based on OpenAI’s recommendations on prompt-engineering (OpenAI 2024) and prior work (Mishra and Chatterjee 2024).
First, prior research showed that models demonstrate improved performance when compelled to reason and justify their decisions (OpenAI 2024; Dammu et al. 2024). Therefore, we mandated GPT-4 to provide direct excerpts from the YouTube video metadata and concise justifications regarding their selected label.
Second, OpenAI (2024) recommended that asking the model to adopt a “persona” in their system can lead to better results from LLMs. Thus, for our task, we prompted GPT-4 to adopt a persona as a “public health expert” with comprehensive knowledge of the COVID-19 pandemic and its misinformation.
Third, we experimented with both zero-shot prompting and few-shot prompting. Zero-shot prompts involve presenting the task to the LLM, only including the label descriptions without accompanying examples or training. Meanwhile, few-shot prompts involve including task-specific examples within the prompt, enabling the pre-trained LLM to condition on the illustrative examples rather than updating its weights. We provided five few-shot examples, each containing video metadata, an assigned label, and reasoning for the label. These few-shot examples were the same as the ones provided to AMT workers on their annotation tasks.
Fourth, we experimented with varying temperature levels to find our task’s most ideal parameter setting. Ranging from 0 to 2, the temperature parameter influences how models generate text. When utilizing lower values for temperature, such as 0, the generated text becomes deterministic, selecting more consistent and coherent outputs.Previous works (Mishra and Chatterjee 2024; Park et al. 2024) also experimented with various temperature values, such as 0.2 and 0.7, for their text annotation tasks and found that lower temperature values (e.g. 0.2) resulted in the best performance. Likewise, we selected relatively lower temperature values: 0, 0.2, and 0.7 for our experiments. We used default settings for other parameters. See Table 12 for our zero-shot prompt and Table 13 for our few-shot prompt.
@@ Appendix H Classifier Performance Results
Performance results for each individual feature and combination of features are presented in Index 0-13 of Table 11. Among the individual features, we observed that video titles yielded the highest performance (0.64 accuracy), suggesting that they are an informative input feature. Similarly, descriptions (0.6 accuracy) and transcripts (0.64 accuracy) demonstrated comparable performance. In contrast to the findings of Medina Serrano, Papakyriakopoulos, and Hegelich (2020), where they identified video comments as highly predictive in detecting COVID-19 misinformation in YouTube videos, our results revealed video comment as the least effective input feature, achieving an accuracy of only 0.49. Combining title, description, and transcript (Index 9) or title, description, transcript, and tags (Index 12) yielded the best performance of 0.69 accuracy. This further indicates that video comments are not informative for our classification task.
@@ Appendix I Significance Testing Details
Appendix K Validations and Robustness Analysis
In this section, we conducted validation checks to reinforce the robustness of our findings.